# Positive Effects of Financing

### v2

@@ Abstract
We show that the classical calibration of the DICE model induces intergenerational inequality: future generations have to bear higher costs relative to GDP. Further, we show that considering stochastic discount rates and stochastic abatement policies, which can be interpreted as successive re-calibration, increases intergenerational inequality (and adds additional risks). Motivated by this, we consider additional financing risks by investigating two modifications of DICE. We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing. To conclude our discussion of options to improve intergenerational equity in an IAM, we propose a modified optimization to keep costs below 3 % of GDP, resulting in more equal distribution of efforts between generations.
@@ 1 Introduction
Climate change is one of the most significant risks of the next centuries [1], as demonstrated by its manifold impacts on society [2]. Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide [3], a transition to (net) zero emissions is necessary to limit global warming to any constant level. A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world. One traditional method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
We focus on the DICE model [6, 7] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [8, 9, 10, 11, 12]. While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points [13, 14, 15, 16, 17, 18, 19], natural feedback processes such as permafrost [9] or abstract catastrophic risks [20, 21]. Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimization function have a strong influence on the resulting optimal pathways [22, 23, 24, 25, 26]. Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks [13] or inequalities in the distribution of damages [27].
@@
Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation. Model extensions updating IAMs [11, 12, 19, 28] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [29] or energy consumption [30], but also general SCC [31] emphasize the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels. Discount rates, in particular, are an essential component of IAM modelling. Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE [7] versus a normative approach to discounting as in the Stern review [22] continues. The assumptions on discounting have already been discussed in the context of intergenerational [32] and intragenerational inequality [33]. Recent studies [34, 35] show that dynamic modelling of discount rates might allow a more precise assessment of the costs of damages and abatement. While our focus lies on the effect of interest rates on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [36], e.g. carbon taxation, [37].
@@
Parallel to this work, an updated DICE-2023 model was developed and published [38] - one of the most prominent updates includes the consideration of uncertainty in the discount rate through the inclusion of consumption growth uncertainty and riskiness of climate investments. Since the final model remains with a deterministic expected discount rate, our contribution remains independent considering extensions with funding mechanisms and a stochastic discount rate model.
@@
We examine the calibration of the DICE model with respect to intergenerational equity. Measuring the costs per GDP of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations. We use an implementation of the DICE model with a flexible time discretization [39] and couple this model with a comprehensive modelling toolbox for financial markets [40] to enable the integration of valuation methods for complex financial products in IAMs. We add stochastic interest rates to the model and introduce a stochastic abatement policy. The stochastic abatement policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calibration of the model. Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that influence intergenerational equity: Financing of abatement costs and non-linear discounting of costs, [41]. As abatement is a planned investment process, we assume that these costs may be financed by loans with stochastic interest rates. Furthermore, as large projects and investments often incur additional costs [42, 43], we introduce increasing financing costs for large payments via a non-linear discounting. These higher costs for high damages can be viewed as (additional) financing risks. We provide an overview of the extensions in fig. 1. More background is provided in the methods section.
@@ 2.1 Intergenerational Inequality
The conditions (7) and (9) are plausible if a single agent optimizes its cost but may be problematic if the costs are distributed between different generations. Note that the two individual contributions in eq. 9 occur at different magnitudes at different times, even different generations. fig. 2 shows the temporal distribution of the weighted functions $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$.
Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$. The calibration is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
Hence, marginal gains of one generation are equated to marginal losses of another generation. The calibration does not equalize the burden but the weighted marginal burden. In this sense, the calibration is indifferent to the temporal distribution of the absolute burden. This may generate intergenerational inequality.
@@ 2.2 Including stochastic interest rates into DICE
We see in Equation 7 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the numeraire ratio $\frac{N(0)}{N(t)}$ (i.e., the discount factor) are important to determine the inter-temporal optimal pathway.
@@
We use a classical Hull-White model - details are provided in the methods (11). The model generates the same expected discount rate as for the deterministic DICE model we compare it with, which implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model. As is well known, from the literature, the level of the discount rate has a large influence on optimal abatement paths - we show in the supplementary information, that an increasing interest rate level defers abatement, while increasing volatility speeds up abatement, as we further discuss in [39].
@@
fig. 3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure, on emissions, optimal abatement, damage and total cost-per-GDP. These quantities are depicted for the classical deterministic DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10th and 90th percentile values.
@@
This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles). It also lowers the abatement and increases the emission.
Such an effect is known under the name of “convexity” in the context of the valuation of financial derivatives: While (10) shows that stochastic interest rates will result in the same cost and emission paths for the deterministic abatement strategy, we do not recover these cost and emissions for the stochastic abatement strategy as cost and emission are convex functions in the interest rates. Note that for a convex function $f$ and a random variable $X$ we have $\mathrm{E}(f(X))\leq f(\mathrm{E}(X))$ (Jensen’s inequality).
@@ 2.3 Including funding of abatement and funding risks of damage into DICE
We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time for interest that has to be paid. Second, we include a non-linear dependency of discounting of the damage cost on the magnitude of the cost. This can be viewed as the additional cost associated with financing risk of very large amounts [41]. Details on this approach and parameters are provided in the methods section 4.3.4.
We find that both modifications improve intergenerational equity in the classical deterministic DICE model, fig. 4, and in the DICE model with stochastic interest rate, fig. 5. In the latter, intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles.
@@
Figure 5: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost. Shaded area shows the 10th to 90th percentiles.
@@
Figures 4 and 5 depict the instantaneous cost-per-GDP ($C(t)/GDP(t)$). For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [48] to calculate burdens experienced during the projected lifetime.
We measure the burden in percentage-of-GDP (section 4.2.3), where we choose 3 % as the neutral level (green). While differences are small for the time covered by population projections (fig. 6 and fig. 7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century. Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset. The deterministic model keeps the 3% limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
@@ 3 Discussion
Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP. This is because the calibration does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change. Absolute cost and its distribution over time are irrelevant to the objective function. This effect is driven by the aggregation of discounted welfare into intergenerational welfare.
Considering stochastic interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost. This indicates that a repeated re-calibration of the DICE model to current interest rate levels will amplify intergenerational inequality.
We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost. Funding of abatement cost may be considered as modelling the backing of abatement costs by government loans. We find that this may have a positive effect on intergenerational equity in terms of the cost distribution as it improves the alignment of abatement and damage cost. The non-linear discounting of damage cost may be considered as modelling financing risk of damage costs, where very large cost have higher financing cost. The non-linear damage cost can be used as a constrain that allows to effectively limit the generational cost to a specific percentage of the GDP.
Using a simple IAM includes several caveats. First, the representation of damages could be improved [49, 11, 12]. Second, the lack of regional resolution compared to, e.g., econometric damage estimates [50, 51] ignores regional inequalities in exposure and potential distributional effects of projected costs [52]. Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways [31, 53].
Modifying the DICE model in a modular way, we address the issue that the classical model does not consider equity between generations. The proposed extensions accounting for the funding cost of abatement and increased financing cost of large damages can improve this aspect of the intertemporal optimization. This is amplified when considering stochastic interest rate risk. Constraining the total costs of mitigation and damages relative to GDP could be used as a simple heuristic to ensure intergenerational equity of climate mitigation.
@@ 4 Methods
The starting point of our investigations and extensions is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research. Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is re-defined as being an Euler-discretisation of a time-continuous model. For the equi-distant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered. Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$. The differences to the classical DICE-2016 are negligible for our work.
@@
Except where otherwise stated, we use the same model components (e.g. abatement cost and damage cost model) as in DICE-2016. A detailed description of our implementation of the DICE model (and its extensions) can be found in [39]. We use the same notation, except that we denote the cost functions with $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
@@
We will recall the calibration procedure first in Section 4.1, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
@@
We then present metrics to analyse the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.2, since the question at which time cost occur is a question of inter-generational equity.
@@
We conclude presenting different extensions to the model, which change the distribution of cost. We discuss these in Section 4.3. We perform our analysis for the classical (deterministic) DICE model, as well as a model extended with stochastic interest rates and a stochastic abatement policy.
4.1 Model Calibration

### v3

@@ Abstract
We show that the optimization procedure of the DICE model induces intergenerational inequality: lacking a mechanism to regulate burden, future generations have to bear higher costs from abatement and damage relative to GDP.
Further, we show that considering uncertainty of discount rates and their feedback to abatement policies, which can be interpreted as successive re-calculation, increases intergenerational inequality (and adds additional risks). Motivated by this, we consider additional financing risks by investigating two modifications of DICE. We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing. To conclude our discussion of options to improve intergenerational equity in an IAM, we propose a modified optimization to keep costs below 3% of GDP, resulting in more equal distribution of efforts between generations.
@@ 1 Introduction
Climate change is one of the most significant risks of the next centuries [1], as demonstrated by its manifold impacts on society [2]. Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide [3], a transition to (net) zero emissions is necessary to limit global warming to any constant level. A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world. One well-established method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
@@
We focus on the DICE model [6, 7] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [8, 9, 10, 11, 12]. While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points [13, 14, 15, 16, 17, 18, 19], natural feedback processes such as permafrost [9] or abstract catastrophic risks [20, 21]. Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimization function have a strong influence on the resulting optimal pathways [22, 23, 24, 25, 26]. Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks [13] or inequalities in the distribution of damages [27]. Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation. Model extensions updating IAMs [11, 12, 19, 28] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [29] or energy consumption [30], but also general SCC [31] emphasize the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels. Discount rates, in particular, are an essential component of IAM modeling. Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE [7] versus a normative approach to discounting as in the Stern review [22] continues. The assumptions on discounting have already been discussed in the context of intergenerational [32] and intragenerational inequality [33]. Recent studies [34, 35] show that dynamic modeling of discount rates might allow a more precise assessment of the costs of damage and abatement. While our focus lies on the effect of uncertainty about discount rates and stochastic adaptation of abatement policies on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [36], e.g. carbon taxation, [37].
@@
Parallel to this work, an updated DICE-2023 model was developed and published [38] - one of the most prominent updates includes the consideration of uncertainty in the discount rate through the inclusion of consumption growth uncertainty and riskiness of climate investments. These uncertainties are translated into certainty-equivalent discount rate, such that the overall model remains deterministic. As we show in eq. 11, even a stochastic interest rate without feedback processes within the model is equivalent to a deterministic approach.
Complementary to the potential uncertainties in future consumption growth and riskiness of climate investments, we consider uncertainty about the rate of time preference $\rho$ itself. For this, we use a standard interest-rate model - the Hull-White model [39] to model stochastic fluctuation of the rate of time preference $\rho$, interacting with an abatement policy adjusting to this uncertainty.
1.1 Extensions to the DICE Model
@@
We examine the optimization of the DICE model with respect to uncertainty in the rate of time-preference, and its consequences for intergenerational equity. Measuring the costs per GDP (section 4.4.5) of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations. We use an implementation of the DICE model with a flexible time discretization [40] and couple this model with a comprehensive modeling toolbox for financial markets [41] to enable the integration of valuation methods for complex financial products in IAMs.
1.2 Interest Rates
We add stochastic interest rates - through modeling the rate of time preference $\rho$ as stochastic - to the model, and introduce a stochastic abatement policy. This policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calculation of the model. Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that influence intergenerational equity: Financing of abatement costs and non-linear discounting of costs, [42]. As abatement is a planned investment process, we assume that these costs may be financed by loans with uncertain interest rates. Furthermore, as large projects and investments often incur additional costs [43, 44], we introduce increasing financing costs for large payments via a non-linear discounting. These higher costs for high damages can be viewed as (additional) financing risks. We provide an overview of the extensions in fig. 1. More background is provided in the methods section.
@@ 2.1 Intergenerational Inequality
The conditions (8) and (10) are plausible if a single agent optimizes its cost but may be problematic if the costs are distributed between different generations. Note that the two individual contributions in eq. 10 occur at different times. fig. 2 shows the temporal distribution of the weighted functions $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$.
Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$. The optimization is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
Hence, marginal gains of one generation are equated to marginal losses of another generation. The optimization does not equalize the burden but the weighted marginal burden. In this sense, the optimization is indifferent to the temporal distribution of the absolute burden. This may generate intergenerational inequality.
@@ 2.2 Including stochastic interest rates into DICE
We see in Equation 8 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the discount factor $\frac{N(0)}{N(t)}$ (i.e., the numeraire ratio) are important to determine the inter-temporal optimal pathway.
@@
We use a classical Hull-White model to model uncertainty in the rate of time preference $\rho$ details are provided in the methods (12). The model generates the same expected discount rate as for the deterministic DICE model. This implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model. As is well known, from the literature, the level of the discount rate has a large influence on optimal abatement paths - we show in the supplementary information, that an increasing interest rate level defers abatement, while increasing volatility speeds up abatement, as we further discuss in [40].
@@
fig. 3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure (VaR), on emissions, optimal abatement, damage and total cost-per-GDP. These quantities are depicted for the classical deterministic DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10th and 90th percentile values.
@@
This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles). This is qualitatively independent of the interest rate level.
Such an effect is known under the name of “convexity” in the context of the valuation of financial derivatives: While (11) shows that stochastic interest rates will result in the same cost and emission paths for the deterministic abatement strategy, we do not recover these cost and emissions for the stochastic abatement strategy as cost and emission are convex functions in the interest rates. Note that for a convex function $f$ and a random variable $X$ Jensen’s inequality implies $f(\mathrm{E}(X))\leq\mathrm{E}(f(X))$.
The strength of the effect increases with increasing interest rate volatility. A corresponding analysis for different volatility levels can be found in Variation of the Interest Rate Volatility of the supplementary information. For the limits cast of vanishing interest rate volatility $\sigma=0$, we recover the classical deterministic DICE model.
@@ 2.3 Including funding of abatement and funding risks of damage into DICE
We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time, resulting in interest that has to be paid. Second, we include a non-linear dependency of the discount factor for the damage cost on the magnitude of the cost. This can be viewed as the additional cost associated with financing risk of very large amounts [42]. Details on this extension and parameters are provided in the methods section 4.4.4.
We find that both extensions improve intergenerational equity in the classical deterministic DICE model, fig. 4, and in the DICE model with stochastic interest rate, fig. 5. In the latter, intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles.
@@
Figure 5: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost. Shaded area shows the 10th to 90th percentiles (VaR).
@@
Figures 4 and 5 depict the instantaneous cost-per-GDP ($C(t)/GDP(t)$). For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [45] to calculate burdens experienced during each birth-year cohorts projected lifetime.
We measure the burden in percentage-of-GDP (section 4.4.5), where we choose 3% as the neutral level (green), [46]. While differences are small for the time covered by population projections (fig. 6 and fig. 7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century. Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset. The deterministic model keeps the 3%-limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
@@ 3 Discussion
Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP. This is because the optimization does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change. Absolute cost and its distribution over time are irrelevant to the objective function. This effect is driven by the aggregation of discounted welfare into intergenerational welfare.
Considering uncertainty in the rate of time-preference and thus overall interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost. This indicates that a repeated re-calculation of the DICE model to current interest rate levels will amplify intergenerational inequality.
We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost. Funding of abatement cost may be considered as modeling the backing of abatement costs by government loans. We find that this may have a positive effect on intergenerational equity in terms of the cost distribution as it improves the alignment of abatement and damage cost. The non-linear discounting of damage cost may be considered as modeling financing risk of damage costs, where very large cost have higher financing cost. The non-linear damage cost can be used as a constrain that allows to effectively limit the generational cost to a specific percentage of the GDP.
Using a simple IAM includes several caveats. First, the representation of damages could be improved [47, 11, 12]. Second, the lack of regional resolution compared to, e.g., econometric damage estimates [48, 49] ignores regional inequalities in interest rate dynamics, exposure, and potential distributional effects of projected costs [50]. Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways [31, 51].
Modifying the DICE model in a modular way, we address the issue that the classical model does not consider uncertainty in interest rates and its effects on equity between generations. The presented extensions accounting for the funding cost of abatement and increased financing cost of large damages reduce intergenerational inequality. This is amplified when considering stochastic interest rate risk due to convexity induced by the DICE objective function. Constraining the total costs of mitigation and damages relative to GDP could be used as a simple objective to ensure intergenerational equity of climate mitigation.
@@ 4 Methods
The starting point of our study is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research.
@@
Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is re-defined as being an Euler-discretisation of a time-continuous model. For the equidistant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered. Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$. The differences to the classical DICE-2016 are negligible for our work.
@@
Except where otherwise stated, we use the same model components (e.g. abatement cost and damage cost model) as in DICE-2016. A detailed description of our implementation of the DICE model and its extensions can be found in [40]. We use the same notation, except that we denote the cost functions by $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
We will first recapitulate the role of the time-preference rate $\rho$ in the DICE model in Section 4.1 as a background to our extension with stochastic time-preference rates.
@@
Second, we recall the optimization procedure in Section 4.2, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
@@
Third, we present metrics to analyze the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.3.
@@
Following upon this concepts, we present our extensions to the model, which change the distribution of cost in Section 4.4. We perform our analysis for the classical (deterministic) DICE model, as well as a model extended with stochastic interest rates and a stochastic abatement policy.
Finally, we introduce the concept of convexity from mathematical finance in section 4.5.
4.1 Time Preference Rate
The DICE model extends a standard neoclassical optimal growth model (the Ramsey model) with economic impacts from climate change.
4.1.1 Recapitulation: Time Preference Rates and the Ramsey Equation
Given that the utility $U$ is a function of the time $t$-consumption $C(t)$, the discounted utility is
$U(C(t))\cdot\exp(-\rho t)\text{.}$
If we assume a declining marginal utility
$\mathrm{d}U(C))\ =\ -\eta\cdot U(C)\cdot\mathrm{d}C\text{,}$
where $\eta$ denotes the elasticity of marginal utility, then we have
$\frac{\mathrm{d}}{\mathrm{d}t}\left(U(C(t))\cdot\exp(-\rho t)\right)$
$\ =\ -\eta\cdot U(C(t))\cdot\frac{\mathrm{d}C(t)}{\mathrm{d}t}\cdot\exp(-\rho t)-\rho\cdot U(C(t))\cdot\exp(-\rho t)$
$\ =\ -(\eta\cdot g+\rho)\cdot U(C(t))\cdot\exp(-\rho t)\ =\ -r\cdot U(C(t))\cdot\exp(-\rho t)\text{.}$
with
$r\ =\ \eta\cdot g+\rho\text{,}$ (1)
where $g=\frac{\mathrm{d}C(t)}{\mathrm{d}t}$ In this context, $r$ is called the social discount rate and Equation (1) is the Ramsey equation.
The (optimal) growth of the consumption $g$ depends on the optimal emission and savings paths, and is an outcome of the model optimization.
The elasticity of marginal utility $\eta$ is an input to the model and usually empirically estimated [52]. For the link between time preference and social discount rate see also [12]. The time-preference $\rho$ is an exogenous parameter and subject to controversial discussions.
That $\rho$ is perceived as exogenous is already transparent from Nordhaus 2017 paper on the DICE model [7], showing a dependency of the social cost of carbon on the chosen (growth-adjusted) discount rate.
In [53] an analysis is performed on the dependency of the social cost of carbon on a distribution of the discount rate. Here, the Bond market interest rate is used as an input to the model.
While [53] considers a distribution of discount rates derived from Bond market rates, there is no feedback effect between the integrated assessment model and the given interest rates.
The novelty of our extensions is to consider a stochastic process of future uncertain discount rates and consider its interaction with an integrated assessment model. This models the possibility that policymakers adapt the abatement strategy to changes in interest rates. Such an adaptation is plausible, because already the static analysis shows that the optimal abatement strategy depends on the level of the interest rate.
We model the time-preference $\rho$ by a classical interest rate model, the Hull-White model. This can be seen as a model for fluctuations in the time-preference, e.g., induced by fluctuation in the financial market. Details are presented in section 4.4.1
4.2 Model Objective
The objective of the DICE model is to find the abatement rate $t\mapsto\mu(t)$ and the savings rate $t\mapsto s(t)$ that maximize a given objective function.